# and the Quantum Hole Argument

### v1

@@ IV Locally Compact Groups and Imperfect Quantum Reference Frames
Let us now see how this group-theoretic formulation of QRFs carries over into our framework. Given $N$ systems carrying the regular representation of $G$ as described above, each with configuration space $X$, our space $\Phi$ of (classical) models is $X^{\times N}$, where $\times$ denotes the Cartesian product. The latter can be partitioned into orbits of the group $G$ (where $G$ acts uniformly, i.e. with the same group element $g\in G$, on each of the $N$ factors in $X^{\times N}$), which can be characterised by the invariant relative configurations of the systems. Any two models that are assigned the same relative configurations are physically equivalent and thus lie in the same orbit. The choice of a subsystem $i$ as a (perfect) QRF can be seen as choosing the section that, for each orbit, picks out the relative configuration $e$ for subsystem $i$. Then, a change from reference frame $i$ to reference frame $j$ is implemented by “reading off” the orbit-dependent group element that relates the new frame $j$ to the old frame $i$ and performing a quantum-controlled transformation on the state. Note that in our framework presented in Sec. II, these group elements are exactly the $g_{\sigma_{i}\to\sigma_{j}}^{(\mathcal{O}_{\varphi})}$ for each orbit $\mathcal{O}_{\varphi}$ The resulting model that acts as a representative for ${\cal O}_{\varphi}$ then lies on the section that assigns the identity element $e$ to the subsystem $j$.
@@
The existence of stabilisers is also discussed in [25] in the context of the space $\Phi$ of models of general relativity. Note that, there, only a subset of the orbits carry the additional symmetry of the stabiliser group, leading to orbits of various “sizes”. As a consequence, the counterpart relation becomes ambiguous. We will discuss the implications of stabilisers in more detail in the context of spacetimes in superposition in Sec. V.
@@
Before closing this Section, let us return to the case of perfect frames and the imperfect ones discussed in the first option above. In this case, a choice of reference frame can be seen as a choice of section, in which the state of the reference system is set to a fixed configuration, in the above case the trivial state |$\ket{e}$, and thereby factorises out. Viewing choices of QRF in this more abstract way (see also discussion in [18]) allows us to embed a large number of earlier works on QRFs in our framework: the translation group which has been our case study presented above [2], the spin rotation group [5, 11], the Lorentz group [20], the conformal symmetries [18], and the quantum diffeomorphisms [19], which will be the subject of Sec. V.
@@ V Quantum Coordinate Fields and Quantum Reference Frame Changes
Finally, let us turn to the example of general relativity with the symmetry group taken to be the diffeomorphism group. This case is of particular interest because of its potential implications for a theory of quantum gravity. Clarifying the role of coordinates, symmetries, and reference frames in classical general relativity was central in its development and is still a subject of active debate in the philosophical literature about various incarnations of the hole argument (see [60, 30] for recent reviews). This clarity is crucial when trying to generalise these ideas to the quantum level, where many of the conceptual difficulties are exacerbated. In this and the following Section, we analyse quantum diffeomorphisms (variants of which have been proposed in [61, 40, 19]) and their conceptual implications by embedding them in the space of models. Moreover, we propose a possible strategy for defining a counterpart relation, i.e. a prescription of what counts as “identity” for points in different spacetimes in superposition, from the fields living thereon.151515 Whenever we use “identity” or “identify” in the following, we do not mean it in the strict sense of items that are utterly identical but rather in the weaker sense of Lewis’ counterpart relation [27, 28, 29] or the “threading” of [25]. Based on this, we will show how to construct a change of QRF for the diffeomorphism group, in the sense of a change of representational convention. This provides the foundation for a new perspective on what one may call a quantum hole argument (cf. [62]) — that is, an extension of the classical hole argument to spacetimes in superposition — which will be explored in the next Section.
@@
The arguments that follow in this Section do not rely on choosing any particular one of these three options. Returning to the formal setup, let us now define a model as the tuple $(\mathcal{M},g_{ab},\chi_{(A)},\tilde{\chi}_{(A)})$, consisting of a manifold, a metric, and two sets of reference fields, whose respective perspectives we are going to adopt below. The set of all these kinematically possible models is the space of models $\Phi$. The diffeomorphism group, $G=\mathrm{Diff}(\mathcal{M})$, partitions it into orbits of physically equivalent models, which can be related to one another by a diffeomorphism. As before, we take the conceptual leap to the quantum realm by considering superpositions of different configurations. This corresponds to a situation in which the metric and the scalar fields have quantum properties in the sense that they can be in different configurations $g_{ab}^{(i)}$, $\chi_{(A)}^{(i)}$, and $\tilde{\chi}^{(i)}_{(A)}$ in each branch $i$ (see Fig. 7). For clarity, we also denote the manifold in each branch by $\mathcal{M}_{i}$, even though all $\mathcal{M}_{i}$ as differentiable manifolds are assumed to be diffeomorphic. In the following, we will focus on the case $i=1,2$; in general, however, we can consider superpositions of an arbitrary number of models.
@@
Regarding the Hilbert space structure of the gravitational field, we want to keep our assumptions minimal. One advantage of the abstract framework of the space of models is that it allows us to forego the technical details of such a structure and focus on the conceptual implications. [^16] In particular, this allows us to go beyond the case of locally compact Lie groups, a restriction that applies to all QRF approaches formulated thus far, but of course not obeyed by the diffeomorphism group. All we need for this framework to apply is that the structure for the quantum description of the gravitational field reflects the symmetry properties of the theory, i.e. can be partitioned into orbits of the diffeomorphism group, and allows for some type of semi-classical states, each peaked around a classical configuration $g_{ab}^{(i)}$ of the metric (see e.g. [64]) and entangled with the corresponding matter field configurations $\chi_{(A)}^{(i)}$ and $\tilde{\chi}_{(A)}^{(i)}$. We would expect such states to arise in the transition from a full theory of quantum gravity to general relativity in a regime where the quantum fluctuations of the gravitational field can be neglected while the coherence of the superposition is still maintained. An example would be the scenario described in Sec. III, in which a massive object in superposition acts as the source for the gravitational field. If the linearity of quantum theory still applies in this context, the spacetime sourced by such a massive object should be in a superposition of macroscopically distinct gravitational fields: one field with the source further to the left, and the other with the source further to the right. In the following, we will further limit our discussion to superpositions of physically inequivalent configurations $(\mathcal{M},g_{ab}^{(i)},\chi_{(A)}^{(i)},\tilde{\chi}_{(A)}^{(i)})$, that is, we will not consider superpositions of models living on the same orbit.
@@ V.1 Reference Frame of $\chi$
In analogy to the example in Sec. III, which we considered from the perspectives of the Earth and the probe particle respectively, we now want to describe a superposition of spacetimes with respect to the $\chi$- or $\tilde{\chi}$-fields respectively. But what does it mean to go into the reference frame associated to, say, the $\chi$-fields precisely? Intuitively, this means using the values of the four scalar fields $\chi_{(A)}$ to label the points and, in particular, “identify” them across the different branches in superposition. Just as we used the Earth or the probe particle to define the zero-point of our one-dimensional system across the different branches, we now use the values of the four scalar fields $\chi_{(A)}$ to label the points of the four-dimensional spacetime manifold. And just as associating the zero-point to a particular particle allowed us to sew together the different branches of the superposition, thus defining a counterpart relation that tells us whether two locations are the same or different across the branches, we can use the values of the $\chi$-fields to identify points across the spacetimes in superposition and thereby represent different field configurations as “inhabiting” the same spacetime points.
@@
To be more precise, we can use the $\chi$-fields to (i) fix a representational convention and (ii) define the associated counterpart relation (see Fig. 8). Firstly, in order to fix a representational convention, we need to pick out one unique model for each orbit. If the four scalar fields $\{\chi_{(A)}\}_{A=0,1,2,3}$ define a bijective map from $\mathcal{M}$ to $\mathbb{R}^{4}$, fixing the values of the fields removes the redundancy induced by the diffeomorphism invariance. [^17] Note that fixing the position of a single particle, as suggested in [33, 34], will not suffice to single out a unique representative for each orbit but rather restricts one to a — still infinite-dimensional — set of possible representatives compatible with the chosen position for the particle. Agreed, given a specific set of scalar fields, especially if we choose to model them more realistically, the $\chi$-fields might repeat their values, so that there is no bijection. In this case, we can still work with an open subregion of $\mathcal{M}$ on which the fields define bijective maps, as is commonly done when defining coordinate fields. Nevertheless, to keep our notation simple, we will continue referring to all of $\mathcal{M}$ in the following. Note further that there are special cases, such as when the $\chi$-fields take on periodic values over time or are homogeneous over a region of space (cf. [65]), for which the non-uniqueness arises from the existence of stabilisers, and thus imperfect QRFs, as discussed at the end of Sec. IV. More precisely, in that case, the redundancy is not lifted entirely and it is not possible to pick a single representative of the orbit. As a consequence, the characterisation and identification of points across a superposition of spacetimes via coincidences of field values become non-unique.
@@
This map associates to each point $p\in\mathcal{M}_{1}$ a unique counterpart $q=\mathrm{C}_{\chi}(p)\in\mathcal{M}_{2}$ and thus provides us with one way of specifying what we mean by “the same” or “different” points across the different spacetimes (see Fig. 9). It thus furnishes a concrete, physically motivated realisation of the counterpart relation of [25]; cf. the discussion of Eq. (2) in Sec. II above. [^20] Alternative realisations include the Kretschmann-Komar coordinates [67, 68], which are based on gravitational degrees of freedom, Geroch’s use of tetrad fields to define limits of spacetimes [39], and Barbour’s treatment of best-matching [69]. In addition, the identification of spacetime points across different spacetimes in superposition has been discussed by Hardy [40] and Jia [41].
More concretely, given a superposition of two models $\varphi_{1}=(\mathcal{M}_{1},g_{ab}^{(1)},\chi_{(A)}^{(1)},\tilde{\chi}_{(A)}^{(1)})$ and $\varphi_{2}=(\mathcal{M}_{2},g_{ab}^{(2)},\chi_{(A)}^{(2)},\tilde{\chi}_{(A)}^{(2)})$, the comparison map $\mathrm{C}_{\chi}$ can be shown to be equivalent to the counterpart relation for the section $\sigma$ on which the $\chi$-fields take the same value $\chi_{\ast}$. Denote by $d^{(1)}$ and $d^{(2)}$ the diffeomorphisms such that $\chi_{\ast}=\chi^{(1)}\circ(d^{(1)})^{-1}=\chi^{(2)}\circ(d^{(2)})^{-1}$. [^21] Note that the action of the diffeomorphism on the fields is the inverse of its action on manifold points — we can see the action of a diffeomorphism $d$ either as moving the points as $p\to d(p)$ or as shifting the fields inversely as $\chi\to d^{-1}_{\ast}(\chi)$. Then, the counterpart relation is given by
@@
Figure 9: Consider two configurations $(\mathcal{M}_{1},g_{ab}^{(1)},\chi_{(A)}^{(1)})$ and $(\mathcal{M}_{2},g_{ab}^{(2)},\chi_{(A)}^{(2)})$. Each point $p\in\mathcal{M}_{1}$ is associated a coordinate value through the field $\chi^{(1)}$. Similarly, $\chi^{(2)}$ assigns a coordinate value to each point $q\in\mathcal{M}_{2}$. The comparison map $\mathrm{C}_{\chi}=(\chi^{(2)})^{-1}\circ\chi^{(1)}$ maps the point $p\in\mathcal{M}_{1}$ to $q\in\mathcal{M}_{2}$ if and only if $\chi^{(1)}(p)=\chi^{(2)}(q)$.
@@ V.2 Changing to the Reference Frame of ${\tilde{\chi}}$
in each branch of the superposition respectively (see Fig. 10(a)).
@@
The second step of the QRF transformation (ii) consists in choosing a new representational convention, which passes through the transformed models, and thereby picking out a new counterpart relation to compare objects across the superposition. This is illustrated in Fig. 10(b). Changing the representational convention does not affect the configuration of the metric and matter fields but it changes the way we compare the spacetimes in superposition. Physically, it means that we now use the $\tilde{\chi}$-fields instead of the $\chi$-fields to label and identify spacetime points across the branches of the superposition. That is, we call two points identical — i.e. we identify them or say that they are counterparts — if and only if the $\tilde{\chi}$-fields take the same value at these points. This is implemented through the new comparison map
@@
(a) Quantum diffeomorphism
(b) Change of representational convention
Figure 10: The two steps making up a QRF transformation for the diffeomorphism group: (i) the application of a quantum diffeomorphism and (ii) a change of representational convention to align with the new models. While the former transforms the models, leaving the section $\sigma$ the same, the latter changes the section from $\sigma$ to $\tilde{\sigma}$ and thereby changes the counterpart relation between any two given models.

### v2

@@ IV Locally Compact Groups and Imperfect Quantum Reference Frames
Let us now see how this group-theoretic formulation of QRFs carries over into our framework. Given $N$ systems carrying the regular representation of $G$ as described above, each with configuration space $X$, our space $\Phi$ of (classical) models is $X^{\times N}$, where $\times$ denotes the Cartesian product. The latter can be partitioned into orbits of the group $G$ (where $G$ acts uniformly, i.e. with the same group element $g\in G$, on each of the $N$ factors in $X^{\times N}$), which can be characterised by the invariant relative configurations of the systems. Any two models that are assigned the same relative configurations are physically equivalent and thus lie in the same orbit. The choice of a subsystem $i$ as a (perfect) QRF can be seen as choosing the section that, for each orbit, picks out the relative configuration $e$ for subsystem $i$. Then, a change from reference frame $i$ to reference frame $j$ is implemented by “reading off” the orbit-dependent group element that relates the new frame $j$ to the old frame $i$ and performing a quantum-controlled transformation on the state. Note that in our framework presented in Sec. II, these group elements are exactly the $g_{\sigma_{i}\to\sigma_{j}}^{(\mathcal{O}_{\varphi})}$ for each orbit $\mathcal{O}_{\varphi}$. The resulting model that acts as a representative for ${\cal O}_{\varphi}$ then lies on the section that assigns the identity element $e$ to the subsystem $j$.
@@
The existence of stabilisers is also discussed in [25] in the context of the space $\Phi$ of models of general relativity. Note that, there, only a subset of the orbits carry the additional symmetry of the stabiliser group, leading to orbits of various “sizes”. [^18] In fact, this means that the space of models can be described as an orbifold (cf. [50]). As a consequence, the counterpart relation becomes ambiguous. We will discuss the implications of stabilisers in more detail in the context of spacetimes in superposition in Sec. V.
@@
Before closing this section, let us return to the case of perfect frames and the imperfect ones discussed in the first option above. In this case, a choice of reference frame can be seen as a choice of section, in which the state of the reference system is set to a fixed configuration, in the above case the trivial state |$\ket{e}$, and thereby factorises out. Viewing choices of QRF in this more abstract way (see also discussion in [18]) allows us to embed a large number of earlier works on QRFs in our framework: the translation group which has been our case study presented above [2], the spin rotation group [5, 11], the Lorentz group [20], the conformal symmetries [18], and the (not locally compact) quantum diffeomorphisms [19], which will be the subject of Sec. V.
@@ V Quantum Coordinate Fields and Quantum Reference Frame Changes
Finally, let us turn to the example of general relativity with the symmetry group taken to be the diffeomorphism group. This case is of particular interest because of its potential implications for a theory of quantum gravity. Clarifying the role of coordinates, symmetries, and reference frames in classical general relativity was central in its development and is still a subject of active debate in the philosophical literature about various incarnations of the hole argument (see [64, 30] for recent reviews). This clarity is crucial when trying to generalise these ideas to the quantum level, where many of the conceptual difficulties are exacerbated. In this and the following section, we analyse quantum diffeomorphisms (variants of which have been proposed in [65, 39, 19]) and their conceptual implications by embedding them in the space of models. Moreover, we propose a possible strategy for defining a counterpart relation, i.e. a prescription of what counts as “identity” for points in different spacetimes in superposition, from the fields living thereon.191919 Whenever we use “identity” or “identify” in the following, we do not mean it in the strict sense of items that are utterly identical but rather in the weaker sense of Lewis’ counterpart relation [27, 28, 29] or the “threading” of [25]. Based on this, we will show how to construct a change of QRF for the diffeomorphism group, in the sense of a change of representational convention. This provides the foundation for a new perspective on what one may call a quantum hole argument (cf. [66]) — that is, an extension of the classical hole argument to spacetimes in superposition — which will be explored in the next section.
@@
The arguments that follow in this section do not rely on choosing any particular one of these three options. Returning to the formal setup, let us now define a model as the tuple $(\mathcal{M},g_{ab},\chi_{(A)},\tilde{\chi}_{(A)})$, consisting of a manifold, a metric, and two sets of reference fields, whose respective perspectives we are going to adopt below. The set of all these kinematically possible models is the space of models $\Phi$. The diffeomorphism group, $G=\mathrm{Diff}(\mathcal{M})$, partitions it into orbits of physically equivalent models, which can be related to one another by a diffeomorphism. As before, we take the conceptual leap to the quantum realm by considering superpositions of different configurations. This corresponds to a situation in which the metric and the scalar fields have quantum properties in the sense that they can be in different configurations $g_{ab}^{(i)}$, $\chi_{(A)}^{(i)}$, and $\tilde{\chi}^{(i)}_{(A)}$ in each branch $i$ (see Fig. 7). For clarity, we also denote the manifold in each branch by $\mathcal{M}_{i}$, even though all $\mathcal{M}_{i}$ as differentiable manifolds are assumed to be diffeomorphic, and are, strictly speaking, part of the theory and not of each individual model. In the following, we will focus on the case $i=1,2$; in general, however, we can consider superpositions of an arbitrary number of models.
@@
Regarding the Hilbert space structure of the gravitational field, we want to keep our assumptions minimal. Let us emphasise that we do not claim to provide a full construction of the Hilbert space for the metric. In particular, we do not resolve the mathematical challenges surrounding the existence of a well-defined measure over the set of all spacetime metrics. While we recognise this as a significant open problem, our focus here lies on the conceptual insights that can be gained in the meantime. What we need for the present framework to apply is that the structure for the quantum description of the gravitational field reflects the symmetry properties of the theory, i.e. can be partitioned into orbits of the diffeomorphism group, and allows for some type of semi-classical states, each peaked around a classical configuration $g_{ab}^{(i)}$ of the metric (see e.g. [68]) and entangled with the corresponding matter field configurations $\chi_{(A)}^{(i)}$ and $\tilde{\chi}_{(A)}^{(i)}$. We would expect such states to arise in the transition from a full theory of quantum gravity to general relativity in a regime where the quantum fluctuations of the gravitational field can be neglected while the coherence of the superposition is still maintained. An example would be the scenario described in Sec. III, in which a massive object in superposition acts as the source for the gravitational field. If the linearity of quantum theory still applies in this context, the spacetime sourced by such a massive object should be in a superposition of macroscopically distinct gravitational fields: one field with the source further to the left, and the other with the source further to the right. From a mathematical point of view, all this implies is that we can consider linear combinations of the aforementioned semi-classical states in the abovementioned Hilbert space for the spacetime metric. In the following, we will further limit our discussion to superpositions of physically inequivalent configurations $(\mathcal{M},g_{ab}^{(i)},\chi_{(A)}^{(i)},\tilde{\chi}_{(A)}^{(i)})$, that is, we will not consider superpositions of models living on the same orbit.
@@ V.1 Reference Frame of $\chi$
In analogy to the example of a massive object in superposition in Sec. III, which we considered from the perspectives of the Earth and the probe particle respectively, we now want to describe a superposition of spacetimes with respect to the $\chi$- or $\tilde{\chi}$-fields respectively. But what does it mean to go into the reference frame associated to, say, the $\chi$-fields precisely? Intuitively, this means using the values of the four scalar fields $\chi_{(A)}$ to label the points and, in particular, “identify” them across the different branches in superposition. Just as we used the Earth or the probe particle to define the zero-point of our one-dimensional system across the different branches, we now use the values of the four scalar fields $\chi_{(A)}$ to label the points of the four-dimensional spacetime manifold. And just as associating the zero-point to a particular particle allowed us to sew together the different branches of the superposition, thus defining a counterpart relation that tells us whether two locations are the same or different across the branches, we can use the values of the $\chi$-fields to identify points across the spacetimes in superposition and thereby represent different field configurations as “inhabiting” the same spacetime points.
@@
To be more precise, we can use the $\chi$-fields to (i) fix a representational convention and (ii) define the associated counterpart relation (see Fig. 8). Firstly, in order to fix a representational convention, we need to pick out one unique model for each orbit. If the four scalar fields $\{\chi_{(A)}\}_{A=0,1,2,3}$ define a bijective map from $\mathcal{M}$ to $\mathbb{R}^{4}$, fixing the values of the fields removes the redundancy induced by the diffeomorphism invariance. [^20] Note that fixing the position of a single particle, as suggested in [33, 34], will not suffice to single out a unique representative for each orbit but rather restricts one to a — still infinite-dimensional — set of possible representatives compatible with the chosen position for the particle. Agreed, given a specific set of scalar fields, especially if we choose to model them more realistically, the $\chi$-fields might repeat their values, so that there is no bijection. In this case, we can still work with an open subregion of $\mathcal{M}$ on which the fields define bijective maps, as is commonly done when defining coordinate fields. Nevertheless, to keep our notation simple, we will continue referring to all of $\mathcal{M}$ in the following. Note further that there are special cases, such as when the $\chi$-fields take on periodic values over time or are homogeneous over a region of space (cf. [wüthrich_2009]), for which the non-uniqueness arises from the existence of stabilisers, and thus imperfect QRFs, as discussed at the end of subsection Sec. IV. More precisely, in that case, the redundancy is not lifted entirely and it is not possible to pick a single representative of the orbit. As a consequence, the characterisation and identification of points across a superposition of spacetimes via coincidences of field values become non-unique.
@@
This map associates to each point $p\in\mathcal{M}_{1}$ a unique counterpart $q=\mathrm{C}_{\chi}(p)\in\mathcal{M}_{2}$ and thus provides us with one way of specifying what we mean by “the same” or “different” points across the different spacetimes (see Fig. 9). It thus furnishes a concrete, physically motivated realisation of the counterpart relation of [25]; cf. the discussion of Eq. (2) in Sec. II above. Alternative realisations include the Kretschmann-Komar coordinates [70, 71], which are based on gravitational degrees of freedom, Geroch’s use of tetrad fields to define limits of spacetimes [40], and Barbour’s treatment of best-matching [72]. In addition, the identification of spacetime points across different spacetimes in superposition has been extensively discussed by Hardy [39] as well as Jia [41].
Returning to the comparison map, given a superposition of two models $\varphi_{1}=(\mathcal{M}_{1},g_{ab}^{(1)},\chi_{(A)}^{(1)},\tilde{\chi}_{(A)}^{(1)})$ and $\varphi_{2}=(\mathcal{M}_{2},g_{ab}^{(2)},\chi_{(A)}^{(2)},\tilde{\chi}_{(A)}^{(2)})$, $\mathrm{C}_{\chi}$ can be shown to be equivalent to the counterpart relation for the section $\sigma$ on which the $\chi$-fields take the same value $\chi_{\ast}$. Denote by $d^{(1)}$ and $d^{(2)}$ the diffeomorphisms such that $\chi_{\ast}=\chi^{(1)}\circ(d^{(1)})^{-1}=\chi^{(2)}\circ(d^{(2)})^{-1}$. [^23] Note that the action of the diffeomorphism on the fields is the inverse of its action on manifold points — we can see the action of a diffeomorphism $d$ either as moving the points as $p\to d(p)$ or as shifting the fields inversely as $\chi\to d^{-1}_{\ast}(\chi)$. Then, the counterpart relation is given by
@@
Figure 9: The comparison map $C_{\chi}$. Consider two configurations $(\mathcal{M}_{1},g_{ab}^{(1)},\chi_{(A)}^{(1)})$ and $(\mathcal{M}_{2},g_{ab}^{(2)},\chi_{(A)}^{(2)})$. Each point $p\in\mathcal{M}_{1}$ is associated a coordinate value through the field $\chi^{(1)}$. Similarly, $\chi^{(2)}$ assigns a coordinate value to each point $q\in\mathcal{M}_{2}$. The comparison map $\mathrm{C}_{\chi}=(\chi^{(2)})^{-1}\circ\chi^{(1)}$ maps the point $p\in\mathcal{M}_{1}$ to $q\in\mathcal{M}_{2}$ if and only if $\chi^{(1)}(p)=\chi^{(2)}(q)$.
@@ V.2 Changing to the Reference Frame of ${\tilde{\chi}}$
in each branch of the superposition respectively (see Fig. 10).
@@
The second step of the QRF transformation (ii) consists in choosing a new representational convention, which passes through the transformed models, and thereby picking out a new counterpart relation to compare objects across the superposition. This is illustrated in Fig. 10. Changing the representational convention does not affect the configuration of the metric and matter fields but it changes the way we compare the spacetimes in superposition. Physically, it means that we now use the $\tilde{\chi}$-fields instead of the $\chi$-fields to label and identify spacetime points across the branches of the superposition. That is, we call two points identical — i.e. we identify them or say that they are counterparts — if and only if the $\tilde{\chi}$-fields take the same value at these points. This is implemented through the new comparison map
@@
Figure 10: A QRF transformation for the diffeomorphism group. It consists of two steps: (a) the application of a quantum diffeomorphism and (b) a change of representational convention to align with the new models. While the former transforms the models, leaving the section $\sigma$ the same, the latter changes the section from $\sigma$ to $\tilde{\sigma}$ and thereby changes the counterpart relation between any two given models.